OpenStudy (anonymous):

Can someone explain to me how to "solve these systems algebraically"? Example: 5x+2y=10 3x+2y=6 I think I understand; I just want to be sure. This is in "Linear Systems" by the way.

OpenStudy (1645323):

#2. y=1.5x+3

OpenStudy (anonymous):

Ok, so that's the answer to the second part? That's all one problem btw.

OpenStudy (jdoe0001):

have you covered system of equations yet?

OpenStudy (jdoe0001):

or is it just 2 separate fucntions:?

OpenStudy (anonymous):

It's a linear system. I'm usually pretty good at this stuff but I' m kind of lost here. lol

OpenStudy (jdoe0001):

ok... lemme try to show it by means of elimination

OpenStudy (jdoe0001):

\(\large \begin{array}{llll} 5x+2y=10&&\quad 5x\cancel{+2y}=10\\ 3x+2y=6&{\color{red}{ \times -1}}\implies &-3x\cancel{-2y}=-6\\ \hline\\ &&\quad 2x\qquad\qquad =4 \end{array}\) so as you can see what "x" is then then you can just substitute that on either equation and solve for "y" to get "y"

OpenStudy (anonymous):

Hmm... Ok, thanks man. :)